OpenStudy (anonymous):

Use the substitution y=2^x to solve for the x in the equation 2^(2x+1)-2^(x+1)+1=2^x

ganeshie8 (ganeshie8):

\(\large 2^{2x+1}-2^{x+1}+1=2^x\) \(\large 2 \times 2^{2x}-2 \times 2^{x}+1=2^x\)

ganeshie8 (ganeshie8):

say \(\large 2^x = t\)

ganeshie8 (ganeshie8):

plug that above, you will see an all familiar quadratic equation, which should be easy to solve :)

OpenStudy (anonymous):

you are a star dude thanks a lot

ganeshie8 (ganeshie8):

aww thnks :) you're wlcme !!